Grade 11Physical Chemistry

The series limit for the Paschen series of hydrogen spectrum occurs at 8205.8A Calculate. (a) Ionization energy of hydrogen atom (b) Wave length of the photon that would remove the electron in the ground state of the hydrogen atom.

Sukesh Seth
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
For a hydrogen molecule, made out of a circling electron bound to a core of one proton, an ionization vitality of2.18 × 10 −18 joule (13.6 electron volts) is required to drive the electron from its most reduced vitality level completely out of the iota. 
λ=8205.8A 
The Rydberg Expression gives: 
1λ=R[1n21−1n22] 
R is the Rydberg Constant = 1.097×10−2nm−1 
n1=2 
n2=4 
These allude to the guideline quantum quantities of the vitality levels which the electrons move from. 
1λ=1.097×10−2[0.25−0.0625]=0.00206nm−1 
λ=485.4nm 
The sub - shells d and p given are unimportant to the question as, in the hydrogen particle, the sub - shells inside a vitality level are deteriorate (of a similar vitality).